OpenStudy (anonymous):

simplify the polynomial 1. 16x^2 + 2 + 2x – 15x^2 – 1 + 14x a. 31x^2 + 16x + 1 b. x^2 – 12x + 1 c. x^2 + 16x + 1 d. X^2 + 16x + 2

OpenStudy (anonymous):

so is it b?

OpenStudy (texaschic101):

Try again...

OpenStudy (anonymous):

d ?

OpenStudy (texaschic101):

your just guessing.....lets work this out together... What is 16x^2 - 15x^2 = ?

OpenStudy (texaschic101):

your wanting to put your like terms together

OpenStudy (texaschic101):

ok....what is 16 - 15 ?

OpenStudy (anonymous):

no i just suck at this haha , but i just tryed to work it out and is it 31x^2 + 16x + 1?

OpenStudy (texaschic101):

no it isn't

OpenStudy (anonymous):

o my god.

OpenStudy (texaschic101):

let me show you......put your like terms together....put your x^2's together, put your x's together and then put your numbers together... 16x^2 - 15x^2 = 1 x^2 which is the same as x^2 now put x's together.. 2x + 14x = 16x now the numbers.. 2 - 1 = 1 Now put down what you have... x^2 + 16x + 1 Do you understand ? If not, ask questions....you need to learn this.....this is really easy

OpenStudy (anonymous):

thanks i understand it alot better now ! can i post the next question and solve it myself and u tell me if its right or not?

OpenStudy (texaschic101):

The x^2's and the x's stay the same......just the numbers change ...also. don't ever add/subtract an x from an x^2. Just like terms.

OpenStudy (texaschic101):

you post the next question and I will check it :)

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

3x^2+6-2x+5x-4x^2+9

OpenStudy (texaschic101):

get your " like " terms together

OpenStudy (texaschic101):

get your x^2's together, then your x's, then your numbers. Make sure you pay attention to your positive and negative signs

OpenStudy (texaschic101):

Let me start you off..... 3x^2 - 4x^2 = - 1x^2 which is the same as -x^2 now you do the x's

OpenStudy (anonymous):

2x - 5x = 3x

OpenStudy (texaschic101):

pay attention to your signs...however the answer you got is right... it should be -2x + 5x = 3x now do the numbers

OpenStudy (anonymous):

thats the only part that confuses me

OpenStudy (texaschic101):

so what did you get for your answer ?

OpenStudy (texaschic101):

oh...your confused by the signs ?

OpenStudy (anonymous):

yeah ^

OpenStudy (texaschic101):

quick lesson.... if the signs are the same, whether they are negative or positive, you just add them and keep the same sign. Example : -2 + (-1) = - 3..........+ 4 + 4 = + 8 or just 8 if the signs are different.....positive and negative, then you subtract and take the sign of the largest number Example : -3 + 6 = + 3 or just 3........-8 + 7 = - 1 These rules do not apply when multiplying or dividing. Just adding and subtracting. Any questions ?

OpenStudy (anonymous):

then what are the rules for multiplying

OpenStudy (texaschic101):

rules for multiplying and dividing are : if the signs are the same, whether negative or positive, then the sign is always positive. Example : -4 x - 4 = 16.....2 x 6 = 12 if the signs are different positive and negative, then the sign will be negative. Example : -4 x 5 = - 20.......3 x (-3) = - 9 These are for multiplying and dividing